OpenStudy (coolcat):

what is m (over) 3 + 1 (over) 4 = m (over) 4. URGENT!

OpenStudy (turingtest):

m/3+1/4=m/4 multiply by 12: 4m+3=3m combine like terms: m=-3

OpenStudy (anonymous):

-3

OpenStudy (coolcat):

thanks